OpenStudy (anonymous):

The pair of points (6, y) and (10, –1) lie on a line with slope 1/4. What is the value of y?

OpenStudy (chriss):

-1=(1/4)(10)+b b=-3.5 y=(1/4)(6)-3.5 y=-2

OpenStudy (anonymous):

chriss please help me